# Yue Wanga and Jonas Berxa‡

### v1

@@ 1 Introduction
The set-up of this paper is as follows. In Section 2 we introduce the chiral Vicsek model and our braiding formalism, showing the different steady-state patterns and illustrating how to map spacetime trajectories into algebraic braids. In Section 3, the braid-topological measures relevant for mixing are introduced an subsequently computed for the different observed patterns. In Section 4, we show a Pareto-optimal front between the writhe and FTBE before individually optimising the FTBE and numerically computing the parameters associated with the optimal mixing pattern. Finally, in Section Conclusions, we conclude and provide avenues for future research.
@@ 2.1 The chiral Vicsek model
Given these interactions, the coupled equations of motion for the particle positions $\mathbf{r_{m}}(t)=(x_{m}(t),\,y_{m}(t))$ and orientations ${\theta_{m}}$, with $m\in\{1,\dots,N\}$, are given by
$\begin{split}{\bf\dot{r}_{m}}&={\bf v}(\theta_{m})+\frac{2\kappa}{\ell}\sum\limits_{r_{mn}<\varepsilon}({\bf r_{m}}-{\bf r_{n}})\,\mathrm{e}^{-\left(r_{mn}/\ell\right)^{2}}-\kappa_{b}\hat{\bf r}_{m}\,H(r_{m}-R)\,,\\ \dot{\theta}_{m}&=\omega-\gamma_{p}\sum\limits_{r_{mn}<\varepsilon}\sin{(\theta_{m}-\theta_{n})}-\gamma_{w}\sin{2(\theta_{m}-\phi_{m}-\frac{\pi}{2})}H(r_{m}-R)+\eta_{m}\,,\end{split}$ (1)
where $H(x)$ is the Heaviside step function and $\eta_{m}$ is Gaussian noise with mean zero and two-point correlation $\langle\eta_{m}(t)\eta_{n}(t^{\prime})\rangle=2D\delta_{mn}\delta(t-t^{\prime})$. For all subsequent simulations we fix $v_{0}=1$, $R=12$, $\kappa=3$, $\kappa_{b}=20$, $\ell=0.3$, $\varepsilon=1$ and $D=0.01$. This particular choice of parameters aligns well with experimental data on Escherichia coli suspensions confined within a microwell 31. The remaining tunable parameters in our system are thus $N,\,\omega,\,\gamma_{p}$ and $\gamma_{w}$.
@@
Fig.  2: The different representative NESS patterns observed in this work. The top row displays patterns in the weak alignment regime for fixed $N=1000$ and increasing $\omega$, while the bottom row illustrates the corresponding patterns in the strong alignment regime for fixed $N=2000$. A colour legend indicating the orientational angle of the active particles is shown in the upper left corner. Abbreviations for each steady-state pattern are given in square brackets, with their full descriptions provided in the main text. For both rows, patterns on the right of the vertical line (OO and CR+O) oscillate periodically between the two side-by-side patterns.
@@ 2.2 Steady-state patterns
In the strong alignment regime (see Fig. 2 bottom row), both CW and CCW edge currents reappear at low values of $\omega$ across all $N$ (not shown). As $\omega$ increases, the system transitions either to a flocked state—MF followed by SF for low $N$, while predominantly SF for moderate $N$—or to a state where a stable spiral (SS) droplet persists in the NESS, for high $N$. For low $N$, increasing $\omega$ reproduces behaviour similar to the weak alignment regime and the systems transitions into the OO phase. For moderate to high $N$, however, the NESS patterns transition from either the SF or SS into a state in which counterrotating currents (CR) emerge; due to the strong boundary alignment dominating over the chirality, CW edge currents emerge, with particles further away from the boundary forming a CCW rotating current. A single-vortex state with a counterrotating boundary layer have been observed in, e.g., confined bacterial suspensions 34. Increasing $\omega$ further leads the particles away from the boundary to form flocks, resulting in a combination of counterrotation and flocking (CR+F), and eventually for very high $\omega$ this leads to counterrotation with oscillation (CR+O), similar to the OO phase, in which a stable spiral is periodically formed and then destroyed near the centre. For moderate $N$, however, this final phase is observed only for very high $\omega\gtrsim 0.6$; for $\omega\lesssim 0.6$ particle oscillations are unstable and the system collapses into travelling bands (TB) that rotate around the system instead.
@@ 2.3 Braiding trajectories
To characterise mixing, we record the spacetime trajectories of a number $n\leq N$ tagged particles and project the trajectories onto a two-dimensional subspace, where one of the dimensions is the temporal coordinate 22. Henceforth, let us assume we project onto the $x-t$ plane. When projecting, particle trajectories can cross at different points in time. Although we projected on the $x-t$ plane, the remaining spatial coordinate $y$ determines whether a single trajectory crosses either behind or in front of others. In this manner the trajectories become entangled, and the entanglement information is encoded in the collection of crossings between particle trajectories. Based on the temporal coordinate, these trajectory crossings can then be sorted into a standard form, where crossings cannot occur at the same time, neglecting the length between crossings. The resulting object is an algebraic braid, which contains all topological information of the particle movement. A visual representation of this process for our model is shown in Fig. 3, where four particles entangle to form a braid.
A crossing of strands $i$ and $i+1$ can be represented by operators $\sigma_{i}^{\pm 1}$, where the positive (negative) exponent indicates an over (under) crossing. These operators can be concatenated into a representation of the braid: $\beta_{n}=\sigma_{i}^{\pm\nu_{i}}\sigma_{j}^{\pm\nu_{j}}...\sigma_{k}^{\pm\nu_{k}}$, where all $\nu=\pm 1$. To convert the trajectories into braids we use the braidlab Matlab package 35, which also allows for the calculation of a myriad of braid measures.
Fig.  3: Mapping of spacetime trajectories of four tagged particles (left) to an algebraic braid (right) by projecting the trajectories on the $x-t$ plane. The resulting braid is $\beta_{4}=\sigma_{2}^{2}\sigma_{1}^{-1}$ with writhe $\rm{Wr}=1$ and braid length $L=3$.
@@ 3.1 The braid writhe
where the angle $\phi$ is between the particle velocity $\mathbf{v}_{i}$ and unit tangent $\mathbf{T}_{i}$ at the position of the particle $i$. The VOP is equal to one when the system exhibits perfect vortex dynamics and equal to zero when there is perfect disorder. It becomes negative when the motion is radial. Edge currents, for instance, have $\rm{VOP}\approx 1$, while oscillatory patterns transition between circular motion $(\rm{VOP}\approx 1)$ and radial motion $(\rm{VOP}<0)$.
@@ 3.2 The finite-time braiding exponent
A more natural way to characterise mixing is to consider material loops surrounding the strands and studying the (exponential) stretching of the loops by applying the braid operators. The growth rate of a loop under the repeated action of a braid then gives the braid entropy. Let us consider a material loop $\ell_{E}$ representing a generating set of the non-oriented fundamental group on the $n$-punctured disk 36, i.e., a loop around “holes” in a disk representing the braid strand starting points, see the left panel of Fig. 5. Acting on the punctures with a braid operator $\sigma_{i}^{\pm 1}$, rotating them in the plane, can lead to the deformation and subsequent stretching of the loop, see the right panel of Fig. 5 for the braid $\beta_{4}=\sigma_{2}^{2}\sigma_{1}^{-1}$. The resulting integral laminations (the set of disjoint nonhomotopic simple closed curves) formed by the action of the braid can be encoded by using Dynnikov coordinates 37, and the relative growth of the number of intersections with the real axis, denoted by the “norm” |.||.|| . |, is a measure for the growth rate of the material lines, and hence the mixing.
The FTBE of a braid $\beta_{n}$ of $n$ particles for an interval $T$ is defined as 23
@@
Fig.  5: The definition of $\ell_{E}$ for a braid of four strands as loops around punctures (coloured circles) that cannot be deformed into a single point. The right panel illustrates the deformation of the loop $\ell_{2}$ (yellow) under the action of the braid $\beta_{4}=\sigma_{2}^{2}\sigma_{1}^{-1}$, for which representative trajectories are shown in Fig. 3. The growth of the number of intersection points between the deformed loop and the real axis (not shown) is a measure of the topological entropy.
@@
Fig.  6: The FTBEn (a,b) and rescaled writhe (c,d) as a function of chirality for $N=1000$ (a,c) and $N=2000$ (b,d) in the weak alignment regime where $\gamma_{p}=0.1,\,\gamma_{w}=1$. The rescaled writhe provides a very good data collapse. Regions of chirality where certain patterns occur are approximately delineated by the vertical dashed lines using Ref. 31 and visual inspection. The patterns are abbreviated as follows: counterclockwise edge current (CCW), multiple flocks (MF), single flock (SF) and orderded oscillation (OO).
[…]
@@ 4 Mixing and Pareto-optimal trade-offs
$\Omega_{n}=-\lambda{\rm Wr}_{n}-(1-\lambda){\rm FTBE}_{n}\,.$ (4)
@@
Doing so results in a parametric front of writhe versus FTBE 43. A concave front would suggest a protocol –a change in the underlying system parameters by tuning $\lambda$– that transitions smoothly from a state maximising the writhe to one maximising FTBE. A convex front, however, suggests that only the two endpoints, i.e., the points with either maximal writhe or FTBE are optimal solutions, given that solutions lying in between are metastable in the sense that they are local minima of equation (4), while the endpoints are global minima. By tuning $\lambda$, a sudden switch from one endpoint to the other can then be achieved, similar to the liquid-gas first-order phase transition in classical thermodynamics. In between, for a linear front, any Pareto-optimal solution becomes optimal at a critical value of the parameter $\lambda$, resembling a first-order phase transition at criticality. Such a linear front implies a proportional trade-off between writhe and FTBE.
@@
To compute the Pareto front, we use a NSGA-II elitist genetic algorithm 44. Fig. 8 shows that the trade-off between the writhe and FTBE for braids with $n=200$ strands is approximately linear, indicating that they are inversely related, confirming our initial suspicions. Increasing the writhe proportionally decreases the maximal mixing efficiency that can be achieved. Due to the computational load involved in numerically determining the Pareto front, however, the accuracy of the results is limited and possibly allows for other front shapes to emerge when accuracy can be improved. Hence, we will not definitively claim here that the front must be linear.
@@
Fig.  8: The Pareto-optimal trade-off (symbols) between the writhe and FTBE for $n=200$. The indicated slope is calculated by a linear fit with goodness-of-fit $R^{2}=0.987$. Patterns associated with the black labelled points are given underneath. Points 2 and 3 present CCW edge currents, while for point 1, which has the highest FTBE, the system evolves into an edge-confined vortex periodically colliding with a moving flock.
@@
To check which steady-state patterns yield the highest mixing efficiency, however, we use the Nelder-Mead simplex method 45 to maximise the FTBEn independently of the writhe. The numerically determined optimal value is ${\rm FTBE}_{200}=0.072$, which is very close to the leftmost endpoint labelled by number 1 in Fig. 8. The parameters corresponding to this state are listed in Table 1, together with a second simulation for $n=50$ tracer particles, confirming the optimal parameter combination to a good degree. The concomitant optimal mixing pattern combines a stable vortex confined at the boundary with the rest of the particles exhibiting ordered oscillations. The system oscillates between the vortex acting as a mixer and subsequently particles being pushed out and new particles pushed in to be mixed. This interaction between the vortex and the rest of the system leads to the observed high degree of mixing.
@@
$n$ $FTBE_{n}$ $N$ $\omega$ $\gamma_{p}$ $\gamma_{w}$
@@
Conclusions

### v2

@@ 1 Introduction
The set-up of this paper is as follows. In Section 2 we introduce the chiral Vicsek model and our braiding formalism, showing the different steady-state patterns and illustrating how to map spacetime trajectories into algebraic braids. In Section 3, the braid-topological measures relevant for mixing are introduced an subsequently computed for the different observed patterns. In Section 4, we show a Pareto-optimal front between the writhe and FTBE before individually optimising the FTBE and numerically computing the parameters associated with the optimal mixing pattern. In Section 5, we compare some of our computed patterns to experimentally observed ones —in both artificial and living systems— and propose a possible experimental system that could be used as a testbed to demonstrate the validity of our results. Finally, in Section 6, we conclude and provide avenues for future research.
@@ 2.1 The chiral Vicsek model
Given these interactions, the coupled equations of motion (EOM) for the particle positions $\mathbf{r_{m}}(t)=(x_{m}(t),\,y_{m}(t))$ and orientations $\theta_{m}$}, with $m\in\{1,\dots,N\}$, are given by
[…]
with ${\bf v}(\theta_{m})=v_{0}(\cos\theta_{m},\sin\theta_{m})$ and where $H(x)$ is the Heaviside step function and $\eta_{m}$ is Gaussian noise with mean zero and two-point correlation $\langle\eta_{m}(t)\eta_{n}(t^{\prime})\rangle=2D\delta_{mn}\delta(t-t^{\prime})$. For all subsequent simulations we fix $v_{0}=1$, $R=12$, $\kappa=3$, $\kappa_{b}=20$, $\ell=0.3$, $\varepsilon=1$ and $D=0.01$. This particular choice of parameters aligns well with experimental data on Escherichia coli suspensions confined within a microwell 31. The remaining tunable parameters in our system are thus $N,\,\omega,\,\gamma_{p}$ and $\gamma_{w}$.
@@
Fig.  2: The different representative NESS patterns observed in this work. The top row displays patterns in the weak alignment regime for fixed $N=1000$ and increasing $\omega$, while the bottom row illustrates the corresponding patterns in the strong alignment regime for fixed $N=2000$. A colour legend indicating the orientational angle of the active particles is shown in the upper left corner. Abbreviations for each steady-state pattern are given in square brackets, with their full descriptions provided in the main text. For both rows, patterns on the right of the vertical line ($O$ and CR+O) oscillate periodically between the two side-by-side patterns.
@@ 2.2 Steady-state patterns
In the strong alignment regime (see Fig. 2 bottom row), both CW and CCW edge currents reappear at low values of $\omega$ across all $N$ (not shown). As $\omega$ increases, the system transitions either to a flocked state—MF followed by SF for low $N$, while predominantly SF for moderate NN—or to a state where a stable spiral (SS) droplet persists in the NESS, for high $N$. For low $N$, increasing $\omega$ reproduces behaviour similar to the weak alignment regime and the systems transitions into the OO phase. For moderate to high $N$, however, the NESS patterns transition from either the SF or SS into a state in which counterrotating currents (CR) emerge; due to the strong boundary alignment dominating over the chirality, CW edge currents emerge, with particles further away from the boundary forming a CCW rotating current. A single-vortex state with a counterrotating boundary layer have been observed in, e.g., confined bacterial suspensions 34. Increasing $\omega$ further leads the particles away from the boundary to form flocks, resulting in a combination of counterrotation and flocking (CR+F), and eventually for very high $\omega$ this leads to counterrotation with oscillation (CR+O), similar to the OO phase, in which a stable spiral is periodically formed and then destroyed near the centre. For moderate $N$, however, this final phase is observed only for very high $\omega\gtrsim 0.6$; for $\omega\lesssim 0.6$ particle oscillations are unstable and the system collapses into travelling bands (TB) that rotate around the system instead.
@@ 2.3 Braiding trajectories
To characterise mixing, we record the spacetime trajectories of a number $n\leq N$ tagged particles and project the trajectories onto a two-dimensional subspace, where one of the dimensions is the temporal coordinate 22. Henceforth, let us assume we project onto the xtx-t plane. When projecting, particle trajectories can cross at different points in time. Although we projected on the xtx-t plane, the remaining spatial coordinate $y$ determines whether a single trajectory crosses either behind or in front of others. In this manner the trajectories become entangled, and the entanglement information is encoded in the collection of crossings between particle trajectories. Based on the temporal coordinate, these trajectory crossings can then be sorted into a standard form, where crossings cannot occur at the same time, neglecting the length between them. Note that a single crossing always involves two adjacent strands in the standard braid form; entangling distant strands necessarily produces additional crossings with intermediate strands due to the projection operation. After each crossing, the strands are relabelled such that the leftmost strand has once again index $1$. The resulting object is an algebraic braid, which contains all topological information of the particle movement. A visual representation of this process for our model is shown in Fig. 3, where four particles entangle to form a braid.
A crossing of strands $i$ and $i+1$ can be represented by operators $\sigma_{i}^{\pm 1}$, where the positive (negative) exponent indicates that strand $i$, counted from the left, crosses under (over) strand $i+1$. These operators can be concatenated into a representation of the braid: $\beta_{n}=\sigma_{i}^{\pm\nu_{i}}\sigma_{j}^{\pm\nu_{j}}...\sigma_{k}^{\pm\nu_{k}}$, where all $\nu=\pm 1$. To convert the trajectories into braids we use the braidlab Matlab package 35, which also allows for the calculation of a myriad of braid measures.
Fig.  3: Mapping of spacetime trajectories of four tagged particles (left) to an algebraic braid (right) by projecting the trajectories on the xtx-t plane. The resulting braid is $\beta_{4}=\sigma_{2}^{2}\sigma_{1}^{-1}$ with writhe $\rm{Wr}=1$ and braid length $L=3$.
@@ 3.1 The braid writhe
where the angle $\phi$ is between the particle velocity $\mathbf{v}_{i}$ and the unit vector $\mathbf{T}_{i}$ in the azimuthal direction, at the position of the particle $i$. The VOP is equal to one when the system exhibits perfect vortex dynamics and equal to zero when there is perfect disorder. It becomes negative when the motion is radial. Edge currents, for instance, have $\rm{VOP}\approx 1$, while oscillatory patterns transition between circular motion $(\rm{VOP}\approx 1)$ and radial motion $(\rm{VOP}<0)$.
@@ 3.2 The finite-time braiding exponent
A more natural way to characterise mixing is to consider material loops surrounding the strands and studying the (exponential) stretching of the loops by applying the braid operators. The growth rate of a loop under the repeated action of a braid then gives the braid entropy. Let us consider a material loop $\ell_{E}$ representing a generating set of the non-oriented fundamental group on the $n$-punctured disk 36, i.e., a loop around “holes” in a disk representing the braid strand starting points, see the left panel of Fig. 5. The loops also encircle an additional point, represented as an empty circle in Fig. 5, which does not participate in the braiding but serves as an anchor for the loops to wrap around. Acting on the punctures with a braid operator $\sigma_{i}^{\pm 1}$, rotating them in the plane, can lead to the deformation and subsequent stretching of the loop, see the right panel of Fig. 5 for the braid $\beta_{4}=\sigma_{2}^{2}\sigma_{1}^{-1}$. The resulting integral laminations (the set of disjoint non-homotopic simple closed curves) formed by the action of the braid can be encoded by using Dynnikov coordinates 37.
The extent to which a sequence of operators folds the material loops can be measured by counting the number of intersections of the loops with the horizontal real axis through the punctures, and tracking how this number changes after applying each operator. This is shown in the right panel of Fig. 5 as black points, the number of which grows after each application of a braid operator. While the figure illustrates this for a single loop $\ell_{2}$, the total number of intersections for all loops $\ell_{E}$ can be counted by the norm $|.|$. In particular, for a braid $\beta_{n}$ consisting of a sequence of operators, the number of intersections of all loops with the real axis is given by $|\beta_{n}\ell_{E}|$. The relative growth of this number is then a measure for the growth rate of the material lines, and hence the mixing.
To characterise this growth, the FTBE of a braid $\beta_{n}$ of $n$ particles for an interval $T$ is defined as 23
@@
Fig.  5: The definition of $\ell_{E}$ for a braid of four strands as loops around punctures (coloured circles) that cannot be deformed into a single point. The rightmost uncoloured point is the base point which does not participate in the braid formed and serves as an anchor for all loops. The right panel illustrates the deformation of the loop $\ell_{2}$ (yellow) under the action of the braid $\beta_{4}=\sigma_{2}^{2}\sigma_{1}^{-1}$, for which representative trajectories are shown in Fig. 3. The growth of the number of intersections (black points) between the deformed loop and the real axis (thin dashed line) is a measure of the growth rate of the material loop.
@@
Fig.  6: The FTBEn (a,b) and rescaled writhe (c,d) as a function of chirality for $N=1000$ (a,c) and $N=2000$ (b,d) in the weak alignment regime where $\gamma_{p}=0.1,\,\gamma_{w}=1$. The rescaled writhe provides a very good data collapse. Regions of chirality where certain patterns occur are approximately delineated by the vertical dashed lines using Ref. 31 and visual inspection. The patterns are abbreviated as follows: counterclockwise edge current (CCW), multiple flocks (MF), single flock (SF) and orderded oscillation ($O$).
Fig.  7: The FTBEn (top) and rescaled writhe (bottom) as a function of chirality for N=500, 2000, 2500N=500,\,2000,\,2500 in the strong alignment regime where $\gamma_{p}=0.5,\,\gamma_{w}=10$. Regions of chirality where certain patterns occur are approximately delineated by the vertical dashed lines using Ref. 31 and visual inspection. The patterns are abbreviated as follows: (counter)clockwise edge current (CCW/CW), multiple flocks (MF), single flock (SF), ordered oscillation ($O$), stable spiral ($S$), traveling bands (TB), counterrotation (CR) plus flocking (+F) or plus oscillations (+O).
@@ 4 Mixing and Pareto-optimal trade-offs
$\Omega_{n}=-\lambda|{\rm Wr}_{n}|-(1-\lambda){\rm FTBE}_{n}\,.$ (4)
@@
Doing so results in a parametric front of absolute writhe versus FTBE 43. A concave front would suggest a protocol –a change in the underlying system parameters by tuning $\lambda$– that transitions smoothly from a state maximising the writhe to one maximising FTBE. A convex front, however, suggests that only the two endpoints, i.e., the points with either maximal writhe or FTBE are optimal solutions, given that solutions lying in between are metastable in the sense that they are local minima of equation (4), while the endpoints are global minima. By tuning $\lambda$, a sudden switch from one endpoint to the other can then be achieved, similar to the liquid-gas first-order phase transition in classical thermodynamics. In between, for a linear front, any Pareto-optimal solution becomes optimal at a critical value of the parameter $\lambda$, resembling a first-order phase transition at criticality. Such a linear front implies a proportional trade-off between writhe and FTBE.
@@
To compute the Pareto front, we use a NSGA-II elitist genetic algorithm 44. Fig. 8 shows that the trade-off between the writhe and FTBE for braids with $n=200$ strands is approximately linear, indicating that they are inversely related, confirming our initial suspicions. Increasing the writhe proportionally decreases the maximal mixing efficiency that can be achieved. Due to the computational load involved in numerically determining the Pareto front, however, the accuracy of the results is limited and possibly allows for other front shapes to emerge when accuracy can be improved. Hence, we will not definitively claim here that the front must be linear. The parameter ranges for the numerical optimisation are chosen as $\omega\in[0,0.6],\,N\in[500,3000],\,\gamma_{p}\in[0.1,$...] and $\gamma_{w}\in[1,$...]; this choice fully covers both the weak and strong alignment regime for $\gamma_{p},\,\gamma_{w}$, as well as the complete $\omega-N$ phase space determined in Ref. 31.
@@
Fig.  8: The Pareto-optimal trade-off (symbols) between the absolute value of the writhe and FTBE for $n=200$. The indicated slope is calculated by a linear fit with goodness-of-fit $R^{2}=0.987$. Patterns associated with the black labelled points are given underneath. Points 2 and 3 present CCW edge currents, while for point 1, which has the highest FTBE, the system evolves into an edge-confined vortex periodically colliding with a moving flock.
@@
To check which steady-state patterns yield the highest mixing efficiency, however, we use the Nelder-Mead simplex method 45 to maximise the FTBEn independently of the writhe, with the same parameter ranges as in the Pareto optimisation. The numerically determined optimal value is ${\rm FTBE}_{200}=0.072$, which is very close to the leftmost endpoint labelled by number 1 in Fig. 8. The parameters corresponding to this state are listed in Table 1, together with a second simulation for $n=50$ tracer particles, confirming the optimal parameter combination to a good degree. The concomitant optimal mixing pattern combines a stable vortex confined at the boundary with the rest of the particles exhibiting ordered oscillations. The system oscillates between the vortex acting as a mixer and subsequently particles being pushed out and new particles pushed in to be mixed. This interaction between the vortex and the rest of the system leads to the observed high degree of mixing.
@@
$n$ $FTBE_{n}$ NN $\omega$ $\gamma_{p}$ $\gamma_{w}$
@@
5 Roadmap to validation
Although our study is numerical, a number of recent experiments display pattern formation strikingly similar to the regimes we identify. For example, Quincke rollers confined in circular microchambers self-organize into steady boundary vortices punctuated by intermittent particle ejections into the interior 46, or into transient flock-like bands 47. When confined to annular geometries containing obstacles, such systems can exhibit an oscillating global chirality 48, similar to our ordered oscillation pattern.
Moving away from artificial to living systems, bacterial monolayers under confinement can develop counter-rotating edge currents 34, suspected to be a generic phenomenon in bacterial suspensions with highly no-slip boundaries, e.g., oil interfaces. Confined to ‘racetrack’-like geometries they behave similarly to Quincke rollers; they form flock-like bands moving unidirectionally along the track 49.
These qualitative parallels suggest that the oscillating vortex regime maximising mixing in our simulations could be realised in bench-top active matter platforms. To guide a concrete realisation, we envision experiments on pear-shaped polystyrene Quincke rollers in a circular microchamber 50 subject to a perpendicular electric field, whose strength determines the particle self-propulsion speed $v_{0}$. The shape anisotropy of such particles imparts an intrinsic chirality while inelastic collisions—both roller-roller and roller-wall—have been shown to lead to the required alignment mechanisms, due to a build-up of spatial correlations 47, 51. A soft, short-range repulsion may arise naturally from combined hydrodynamic and induced dipolar interactions 52. Together, these mechanisms account for every term in the equations of motion (1), demonstrating that our model is fully amenable to experimental validation and that the optimised mixing phase can, in principle, be observed.
6 Conclusions